# SIGNALS on the mixing of oxygen and nitrogen in the spiral galaxy NGC 6946

### v1

@@ 5 Homogeneity scales of oxygen and nitrogen
Our conclusion is supported by the results of recent high-resolution simulations of a Milky Way-type galaxy presented by Zhang et al. (2024), who tracked seven different chemical elements, including O and N, following their injection into the ISM on a star-by-star basis. Zhang et al. (2024) find that all the elements they considered are well correlated with each other (C representing a mild exception), and that N has virtually the same spatial statistics of the s-process elements Ba and Ce, generated by AGB stars, and deviating only marginally from that of O, S and Mg, released by core-collapse supernovae. N is found to have a slightly longer correlation length than O. The much larger injection width for N (and other AGB-related species) compared to the KT18 model is justified by the longer lifetimes of intermediate-mass stars, which implies longer timescales for the release of their nucleosynthetic products as well as a drift from their birth sites. This is in overall agreement with our findings and arguments, even though we measure a slightly smaller correlation for N compared to O, rather than the opposite effect as found by Zhang et al. (2024).
As a test, we repeated the analysis included in this section using chemical abundances calculated from the R calibration method proposed by Pilyugin & Grebel (2016), although this violates our initial requirement of avoiding diagnostics that combine nitrogen and oxygen lines (Section 3). In this case we find that the correlation curves for the two elements are virtually indistinguishable. However, since the oxygen and nitrogen abundances derived from the R calibration are highly correlated, possibly in an artificial way (Section 3.1), we argue that this result could also be interpreted as a spurious effect in the absence of an alternative abundance determination technique, as used here.
@@ Acknowledgments
Based on observations obtained with SITELLE, a joint project between Université Laval, ABB-Bomem, Université de Montréal and the CFHT with funding support from the Canada Foundation for Innovation (CFI), the National Sciences and Engineering Research Council of Canada (NSERC), Fond de Recherche du Québec - Nature et Technologies (FRQNT) and CFHT. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the indigenous Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.

### v2

@@ 5.2 Semivariogram
The latter conclusion is supported by the results of recent high-resolution simulations of a Milky Way-type galaxy presented by Zhang et al. (2024), who tracked seven different chemical elements, including O and N, following their injection into the ISM on a star-by-star basis. Zhang et al. (2024) find that all the elements they considered are well correlated with each other (C representing a mild exception), and that N has virtually the same spatial statistics of the s-process elements Ba and Ce, generated by AGB stars, and deviating only marginally from that of O, S and Mg, released by core-collapse supernovae. N is found to have a slightly longer correlation length than O. The much larger injection width for N (and other AGB-related species) compared to the KT18 model is justified by the longer lifetimes of intermediate-mass stars, which implies longer timescales for the release of their nucleosynthetic products as well as a drift from their birth sites. This is in overall agreement with our findings and arguments, even though we measure a slightly smaller correlation for N compared to O, rather than the opposite effect as found by Zhang et al. (2024).
@@ Acknowledgments
Based on observations obtained with SITELLE, a joint project between Université Laval, ABB-Bomem, Université de Montréal and the CFHT with funding support from the Canada Foundation for Innovation (CFI), the National Sciences and Engineering Research Council of Canada (NSERC), Fond de Recherche du Québec - Nature et Technologies (FRQNT) and CFHT. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the indigenous Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. We thank Jefeng Li for his comments on our calculation of the two-point correlation function in an earlier draft.